# ToolACE-R: Tool Learning with Adaptive Self-Refinement

### v1

@@ 3.3 Adaptive Self-Refine Iterative Inference
We propose an adaptive self-refine procedure that enables ToolACE-R to iteratively refine its own output and autonomously determine when to halt the iteration. The details are introduced in Algorithm 1. Specifically, given a user query $q$ and candidate tools $T$, ToolACE-R first generates the tool invocation $A_{0}$. In each subsequent iteration, ToolACE-R leverages the answer from the previous iteration as context and refines it. As mentioned earlier, ToolACE-R has learned that when the answer is sufficiently accurate, it can produce the same answer in subsequent refinement. Therefore, the iteration continues until the answers from two consecutive iterations are identical. We refer to this procedure as adaptive self-refine, where the model itself determines when to stop the iterations. To prevent infinite iteration, we set a maximum iteration limit $n$.
@@
0:  Model $\theta$, query $q$, tools $T$, refine prompt $r$, time $n$
0:  Refined final answer $A$
1:  $A_{0}=f_{\theta}(\langle q,T\rangle)$  // First get the direct answer $A_{0}$
@@
3:     $A_{i}=f_{\theta}(\langle q,T\rangle,A_{i-1},r)$  // Get the refined answer
@@
8:  return  $A_{n}$  // $n$ is the maximum iteration time
Through adaptive self-refine, ToolACE-R is able to improve its performance, particularly when facing difficult cases.
4 Experiments
Non-Live Live Overall
Models Simple Multiple Parallel
Parallel
Multiple
Simple Multiple Parallel
Parallel
Multiple
Non-live Live Overall
GPT-4o-2024-11-20 79.42 95.50 94.00 83.50 84.88 79.77 87.50 75.00 88.10 80.75 84.43
Gemini-1.5-Pro-002 78.33 93.50 92.5 90.00 82.95 77.40 87.50 83.33 88.58 78.68 83.63
GPT-3.5-Turbo-0125 77.92 93.50 67.00 53.00 80.62 78.63 75.00 58.33 72.85 78.61 75.73
DeepSeek-V3 78.67 95.5 91.00 91.50 83.72 82.15 81.25 62.50 89.17 82.09 85.63
Hammer2.1-7B (FC) 78.08 95.00 93.50 88.00 76.74 77.4 81.25 70.83 88.65 77.20 82.92
xLAM-7B-r (FC) 74.25 95.50 81.00 73.50 72.09 74.93 50.00 62.50 81.06 73.87 77.47
ToolACE-R (FC) 79.92 96.50 93.50 93.00 81.78 81.01 87.50 83.33 90.73 81.27 86.00
  + Adaptive Self-Refine 81.17 96.00 94.00 93.00 82.95 81.58 87.50 83.33 91.04 81.94 86.49
Table 1: Accuracy comparison on BFCL (Last updated on 2025-01-18). The best results in each category are marked in bold. The second best results are underlined. FC indicates the models are fine-tuned for function calling.
4.1 Experimental Settings
Datasets and Models.
We leverage a subset of ToolACE Liu et al. (2024b) training data as our original full training dataset. The dataset contains multiple types of tool calling data, including cases with multiple candidate tools and parallel function callings, in both Chinese and English. For base model to be trained, we select LLaMA3.1-8B-Instruct AI@Meta (2024) for our main experiments. We also conduct experiments on other base models, including Qwen2.5-Instruct-series Yang et al. (2024) (0.5B, 1.5B, 3B and 7B) and Mistral-7B-Instruct-v0.2 Jiang et al. (2023), to validate the generalizability of ToolACE-R. We compare with the state-of-the-art API-based models, including GPT-series [^1] https://chatgpt.com and Gemini-1.5-Pro [^2] https://gemini.google.com/, as well as open-source models like DeepSeek-V3 Liu et al. (2024a), Llama3.1-8B/70B-Instruct AI@Meta (2024) and Qwen2.5-7B-Instruct Yang et al. (2024). We also compare with fine-tuned tool calling models like ToolACE-8B Liu et al. (2024b), Hammer2.1-7B Lin et al. (2024) and xLAM-7b-r Zhang et al. (2024).
Benchmarks and Evaluation.
We conduct experiments on several representative benchmarks, including the Berkeley Function Call Leaderboard (BFCL) Yan et al. (2024), ACEBench Chen et al. (2025), API-Bank Li et al. (2023) and ToolAlpaca Tang et al. (2023). To fully explore the model’s potential while minimizing external interference and simplifying data preparation, we mainly evaluate performance on single-turn tool calling queries. For BFCL, we evaluate both the Non-live and Live subsets, which correspond to synthetic test cases and real-world scenarios, respectively. Each subset includes four categories: Simple, Multiple, Parallel, and Parallel Multiple. Simple and Multiple cases both involve a single invoked tool, but in Multiple cases, there are multiple candidate tools. Parallel and Parallel Multiple cases require invoking multiple tools simultaneously. For ACEBench, we evaluate only the English normal category, excluding multi-turn cases. This includes the test samples in atom category and single-turn category. For API-Bank and ToolAlpaca, we focus on the first step tool calling samples, disregarding further tool or retrieval feedback. Performance is reported in terms of accuracy for all benchmarks.
Implementation Details.
Given resource constraints, we employ the parameter-efficient fine-tuning method LoRA Hu et al. (2022). All model modules are configured for LoRA fine-tuning, with a rank of 16 and an alpha value of 32. Training is performed with a global batch size of 64 and a learning rate of $1\times 10^{-4}$, following a cosine learning rate schedule with a warmup ratio of 0.1. For data selection, we set the generation temperature to 1.0 to encourage diversity and allow the model to explore a broader range of outputs. Each sample is generated up to 8 times (pass@8) to ensure potential. For the constructed refined data, we apply a loss mask to the first assistant turn of the conversation, as it may involve an incorrect tool invocation, making it unsuitable for the model to learn from this turn. During evaluation, we adopt greedy search to ensure stability and impose a maximum iteration limit of $n=5$.
4.2 Main Results
Results on BFCL.
  • The performance gap between API-based and open-source models is minimal. Large open-source models like DeepSeek-V3 and Llama3.1-70B-Inst show comparable or even superior overall performance to GPT-4o and Gemini-1.5-Pro.
  • Our model, ToolACE-R which leverages our proposed iterative training procedure, achieves the best overall performance, surpassing both large API-based and open-source models.
  • The adaptive self-refinement mechanism further enhances the performance of ToolACE-R, with larger improvement observed in more complex categories (Live), highlighting the effectiveness of our inference method.
Models ACEBench APIBank ToolAlpaca
GPT-4o-2024-11-20 87.00 77.16 83.87
GPT-4o-mini-2024-07-18 81.63 73.10 83.87
Llama3.1-8B-Inst 50.63 60.41 79.03
ToolACE-R 83.88 75.89 85.48
  + Adaptive Sef-Refine 84.00 81.22 88.71
Table 2: Accuracy comparison on ACEBench, API-Bank and ToolAlpaca benchmarks.
Results on More Benchmarks.
Models Non-live Live Overall
Llama3.1-8B-Inst (Base) 84.37 69.28 76.83
ToolACE-R (Full) 91.04 81.94 86.49
  - Adaptive Self-Refine 90.73 81.27 86.00
    - Iterative Training 88.12 80.24 84.18
      - Data Selection 88.96 81.13 85.04
        - Refine Data 88.75 80.09 84.42
Table 3: Ablation Study of our proposed modules.
Ablation Study.
To investigate the impact of the proposed modules, we conduct an ablation study by progressively removing them. The corresponding evaluation results on the BFCL dataset are shown in Table 3. As observed, removing both adaptive self-refinement and iterative training leads to a noticeable performance degradation. However, further removing data selection does not result in a continued drop in performance. In fact, performance improves. This suggests that model-aware data selection should be combined with iterative training to fully leverage the entire training set, as the initial data selection (performed by the base model) typically only captures a small subset of the available data (mostly easy ones). Iterative selection is needed to capture more complex cases, as the trained model is improved. Finally, removing the constructed refine data also negatively impacts performance, validating that self-refine data acts as a form of data augmentation, also enhancing the model’s direct output.
4.3 Effects of Iterative
In this subsection, we examine the effects of our two iterative modules, i.e., model-aware iterative training and adaptive self-refine iterative inference.
Figure 2: Performance of each iteration during training. “Amount” indicates the amount of training samples in that iteration with our data selection module. Iteration 0 means the model before training, i.e., Llama3.1-8B-Inst.
Iterative Training.
Figure 2 illustrates the results of the iterative training process, showing both the accuracy performance and the amount of training data selected at each iteration. Iteration 0 corresponds to the base model, Llama3.1-8B-Inst. As depicted, the quantity of selected samples increases monotonically across iterations, suggesting that the model’s performance on the training set improves, as the selected samples are those the model is more likely to answer correctly (as indicated by pass@k). On the other hand, accuracy improves rapidly during the first two iterations, reaching its peak at iteration 3. Subsequently, accuracy plateaus, with no further improvement in later iterations. This suggests that excessive training could lead to overfitting on the training set.
Iterative Inference.
To assess the effects of our adaptive self-refine method for iterative inference, we conducted additional experiments comparing it with two other inference-time scaling approaches. The first, termed "Vanilla Self-Refine," always selects the last answer at each iteration as the final answer. The second method, "Self-Consistency at Itr n", chooses the majority answer from the answers collected during iterations 1 to n. The results of these inference methods on BFCL, all based on the same model, are shown in Figure 3.
As observed, the performance of "Vanilla Self-Refine" fluctuates across iterations, indicating that the model does not always refine the answer correctly. In contrast, our "Adaptive Self-Refine" method demonstrates more stability across iterations, suggesting that adjacent identical answers serve as a reliable signal for stopping further iterations. The "Self-Consistency at Itr n" method likely represents an upper bound for answer selection, with our method achieving competitive performance compared to this strong baseline. Furthermore, the performance of our method remains consistent across different iteration counts (3, 5, 10), likely because most cases terminate early, well before reaching the maximum iteration limit. This observation is supported by the average iteration time of 2.4, as indicated by our statistics on BFCL.
4.4 Effects of Data Selection
In this subsection, we investigate the effects of our model-aware data selection module. First, we explore the impact of the $k$ value when using pass@k for sample selection. Next, we examine how the use of data selected by different models affects model performance, highlighting the impact of model-aware selection.
Models Direct Refine Amount
Select with Pass@8
  First Iteration 80.24 80.61 120K
  Best Iteration 81.27 81.94 150K
Select with Pass@1
  First Iteration 78.90 78.53 94K
  Best Iteration 79.57 80.01 143K
Table 4: Accuracy performance on BFCL Live category when data selection with Pass@8 and Pass@1. “Direct” and “Refine” indicate the results with direct output and adaptive self-refine, respectively. “Amount” corresponds to the amount of selected training samples.
Data Selection with Pass@k.
In our data selection criterion, we retain training samples that the base model can answer (approximately) correctly within at most $k$ tries (pass@k). In this section, we examine the effects of varying $k$ values. Specifically, we report the evaluation results using pass@8 and pass@1 in Table 4. As shown, although the final number of selected samples is similar between the two methods (150K vs. 143K), the accuracy performance for pass@8 is considerably higher. The performance gap is already pronounced in the first iteration, suggesting that a more lenient selection criterion plays a crucial role in maximizing model potential and shaping the model’s distribution. Notably, this effect does not diminish over additional iterations. Based on these findings, we select a larger $k$ value for our experiments.
Model-aware Selection.
In our data selection process, we propose selecting training samples in a model-aware manner, ensuring that a sample is chosen only when the base model can answer it (almost) correctly. This is based on the assumption that samples aligning with the model’s distribution are more suitable for training. To further assess the impact of such a model-aware approach, we conduct an additional experiment. Instead of using data selected by the base model Llama3.1-8B-Inst, we train Llama3.1-8B-Inst with data selected by Qwen2.5-3B-Inst and Qwen2.5-7B-Inst, as well as with the intersection of all three models’ selected data – i.e., only the samples that all three models perform well on are included in the dataset. The accuracy results of the trained models, along with the corresponding training data amounts, are presented in Table 5. As shown, performance drops when using data selected by other models, even when the data volume is larger. The intersection dataset, which contains less data, achieves comparable results. This observation underscores the benefits of model-aware data selection.
Models Accuracy Amount
ToolACE-R 80.61 120K
Data selected with Qwen2.5-3B 79.50 125K
Data selected with Qwen2.5-7B 80.38 134K
Intersection of Three Datasets 80.27 108K
Table 5: Accuracy performance on BFCL Live category in the first interation when using different selected data to train Llama3.1-8B-Inst, along with the data amount. “Qwen2.5-3/7B” is short for Qwen2.5-3/7B-Inst.
4.5 Further analysis
Models Non-live Live Overall
Raw (Qwen2.5-7B-Inst) 84.38 72.17 78.28
  + Direct Data 88.29 78.39 83.34
  + ToolACE-R 89.23 81.20 85.21
Raw (Mistral-7B-Inst-v0.2) 55.25 55.07 55.16
  + Direct Data 84.06 70.17 77.12
  + ToolACE-R 87.98 80.01 84.00
Table 6: Accuracy performance of Different backbones on BFCL. “+ Direct Data” indicates the model directly trained with the full original set, while “+ ToolACE-R” is the result with our method, both iterative training and adaptive self-refine are applied.
Different Backbones.
Model Size Scaling.
Figure 4: Performance of models in Qwen2.5-Series. “Base” refers to the Qwen2.5-Instruct models, while “Direct” and “Adaptive Self-Refine” represent the results of models trained using our method and evaluated with direct output and Adaptive Self-Refine, respectively.
We further examine whether our method can be applied to models of varying sizes and whether performance improves as model size increases. For this, we use Qwen2.5-Instruct models of different sizes, including 0.5B, 1.5B, 3B, and 7B, as base models. The results, shown in Figure 4, display both the original performance and the performance after applying ToolACE-R. The direct output of ToolACE-R consistently improves performance across all sizes, particularly for smaller models like 0.5B and 1.5B, and the Adaptive Self-Refine inference technique further enhances these results. This demonstrates the effectiveness of ToolACE-R across a range of model sizes.
5 Conclusion
In this paper, we introduce ToolACE-R, a tool learning technique designed to enable large language models (LLMs) to effectively leverage external tools. ToolACE-R consists of two key components: model-aware iterative training and adaptive self-refine iterative inference. Model-aware iterative training maximizes the model’s potential by iteratively selecting the most appropriate samples for training. Adaptive self-refine further enhances performance by allowing the model to autonomously determine when to stop refining its own answers during iterative inference. Experimental results demonstrate that ToolACE-R consistently improves performance across various tool-calling benchmarks and is effective with different backbone models. Additionally, our analyses highlight the impact of the proposed modules on model performance.
Limitations
We summarize the limitations in the following two points:
First, due to resource constraints, we conduct experiments only on models with up to 8B parameters, using LORA. As shown in the scaling performance results in Figure 4, we have not yet observed any bottlenecks preventing larger models from benefiting from our ToolACE-R method.
Second, our focus is on single-turn tool-calling scenarios, excluding the multi-turn and multi-step cases for simplicity. However, we believe that single-turn cases are a fundamental aspect of tool learning. The significant improvements observed across several representative benchmarks demonstrate the effectiveness of our method. We plan to explore ToolACE-R in more complex scenarios in future work.
Appendix A Appendix
A.1 Detailed Introduction to Benchmarks

### v2

@@ 3.3 Adaptive Self-Refine Iterative Inference
We propose an adaptive self-refine procedure that enables ToolACE-R to iteratively refine its own output and autonomously determine when to halt the iteration. Details are shown in Figure 1 (right hand side) and Algorithm 1 in Appendix. Specifically, given a user query $q$ and candidate tools $T$, ToolACE-R first generates the tool invocation $A^{\prime}$. In each subsequent iteration, ToolACE-R leverages the answer from the previous iteration as context and refines it. As mentioned earlier, ToolACE-R has learned that when the answer is sufficiently accurate, it can produce the same answer in subsequent refinement. Therefore, the iteration continues until the answers from two consecutive iterations are identical. We refer to this procedure as adaptive self-refine, where the model itself determines when to stop the iterations. To prevent infinite iteration, we set a maximum iteration limit $n$. Through adaptive self-refine, ToolACE-R is able to improve its performance, particularly when facing difficult cases.
4 Experiments
4.1 Experimental Settings
Datasets and Models.
We leverage a subset of ToolACE [19] training data as our original training data. The dataset contains multiple types of tool calling data, including cases with multiple candidate tools and parallel function callings, in both Chinese and English. We select LLaMA3.1-8B-Instruct [2] for base model in our main experiments. We also conduct experiments on other base models, including Qwen2.5-Instruct-series [39] (0.5B, 1.5B, 3B and 7B) and Mistral-7B-Instruct-v0.2 [11], to validate the generalizability of ToolACE-R. We compare with the state-of-the-art API-based models, including GPT-series [^1] https://chatgpt.com and Gemini-2.5-Pro [^2] https://gemini.google.com/, as well as open-source models like DeepSeek-V3 [17], Llama3.1-8B/70B-Instruct [2] and Qwen2.5-7B-Instruct [39]. We also compare with fine-tuned tool calling models like ToolACE-8B [19], Hammer2.1-7B [16] and xLAM-2-8b-fc-r [41, 24].
Benchmarks and Evaluation.
We conduct experiments on several representative benchmarks, including the Berkeley Function Call Leaderboard (BFCL) [38], ACEBench [5], API-Bank [15] and ToolAlpaca [33]. To minimize external interference and simplifying data preparation, we mainly evaluate performance on single-turn tool calling queries. For BFCL, we evaluate both the Non-live and Live subsets, which correspond to synthetic test cases and real-world scenarios, respectively. Each subset includes four categories: Simple, Multiple, Parallel, and Multiple Parallel. For ACEBench, we evaluate only the English normal category, excluding multi-turn cases. This includes the test samples in atom category and single-turn category. For API-Bank and ToolAlpaca, we focus on the first step tool calling samples, disregarding further tool or retrieval feedback. All performance is reported in terms of accuracy, where only cases with entirely correct tool calling are counted as correct.
Implementation Details.
Given resource constraints, we employ the parameter-efficient fine-tuning method LoRA [9]. All model modules are configured for LoRA fine-tuning, with a rank of 16 and an alpha value of 32. Training is performed with a global batch size of 64 and a learning rate of $1\times 10^{-4}$, following a cosine learning rate schedule with a warmup ratio of 0.1. For difficulty estimation, we set the generation temperature to 1.0 to encourage diversity and allow the model to explore a broader range of outputs. Model generates 8 times (i.e. $k=8$) for each sample to ensure robustness. During data selection, we set $\alpha=0$ and $\beta=0.9$, which means that we discard samples that are extremely easy (difficulty score equals to 0, i.e., the model always answer fully correctly) or too difficult (most attempts are total incorrect). For the constructed refinement data, we apply a loss mask to the first assistant turn of the conversation, as it may involve an incorrect tool invocation, making it unsuitable for the model to learn from this turn. During evaluation, we adopt greedy search to ensure stability. Unless otherwise specified, , the maximum iteration time for adaptive self-refine is set to $n=5$ .
4.2 Main Results
Non-Live Live Overall
Models Simple Multiple Parallel
Multiple
Parallel
Simple Multiple Parallel
Multiple
Parallel
Non-live Live Overall
GPT-4o-2024-11-20 77.17 95.00 93.50 85.00 84.50 79.30 87.50 70.83 87.67 80.24 83.96
GPT-4.1-2025-04-14 80.50 94.00 93.00 87.50 85.66 76.54 93.75 75.00 88.75 78.46 83.60
Gemini-2.5-Pro-Exp-03-25 67.50 91.00 83.50 73.00 79.46 64.96 81.25 79.17 78.75 68.17 73.46
DeepSeek-V3 78.67 95.50 91.00 91.50 83.72 82.15 81.25 62.50 89.17 82.09 85.63
Hammer2.1-7B (FC) 78.08 95.00 93.50 88.00 76.74 77.40 81.25 70.83 88.65 77.20 82.92
xLAM-2-8b-fc-r (FC) 73.08 93.50 87.00 84.00 74.81 66.29 56.25 50.00 84.40 67.51 75.95
ToolACE-R (FC) 79.75 97.00 95.50 91.50 82.95 81.10 93.75 87.50 90.94 81.72 86.33
Table 1: Accuracy comparison on BFCL (Last updated on 2025-04-25). The best results in each category are marked in bold. The second best results are underlined. FC indicates the models are fine-tuned for function calling.
Results on BFCL.
  • The performance gap between API-based and open-source models is minimal. Large open-source models like DeepSeek-V3 and Llama3.1-70B-Inst show comparable or even superior overall performance to GPT models and Gemini-2.5-Pro.
  • Our model, ToolACE-R, which leverages our proposed model-aware iterative training procedure, achieves the best overall performance, surpassing both large API-based and open-source models.
To further assess the effectiveness of our adaptive self-refine method for iterative inference, we conducted experiments comparing with two other inference-time scaling approaches. The first, termed "Vanilla Self-Refine," always selects the last answer at each iteration as the final answer. The second method, "Self-Consistency at Itr n", chooses the majority answer from the answers collected during iterations 1 to n. The results of these inference methods on BFCL, are shown in Figure 2.
As observed, the performance of "Vanilla Self-Refine" fluctuates across iterations, indicating that the model does not always refine the answer correctly. In contrast, our "Adaptive Self-Refine" method demonstrates more stability across iterations, suggesting that adjacent identical answers serve as a reliable signal for stopping further iterations. Our statistics also show that the average iteration time of our method on BFCL is $2.4$ (when maximum iteration time $n=5$), indicating that most cases terminate early thus achieving consistent performance. Our method achieves competitive performance compared to the strong baseline "Self-Consistency at Itr 10", further validating advantages.
Models ACEBench APIBank ToolAlpaca
GPT-4o 87.00 77.16 83.87
GPT-4o-mini 81.63 73.10 83.87
Llama3.1-8B-Inst 50.63 60.41 79.03
ToolACE-R 83.88 75.89 85.48
  + Adaptive SR 84.00 81.22 88.71
Table 2: Accuracy comparison on ACEBench, API-Bank and ToolAlpaca benchmarks. “SR” is short for “Self-Refine”.
Models Non-live Live Overall
ToolACE-R + Adaptive SR 91.32 82.34 86.83
ToolACE-R 90.94 81.72 86.33
  - Data Selection 88.96 81.13 85.04
    - SR Data 88.75 80.09 84.42
Llama3.1-8B-Inst (base) 84.37 69.28 76.83
Table 3: Ablation Study of our proposed modules on BFCL. “SR” is short for “Self-Refine”. The modules are progressively removed.
Results on More Benchmarks.
Ablation Study.
To evaluate the contribution of each proposed module, we conduct an ablation study by progressively removing them. The corresponding results on the BFCL dataset are presented in Table 3. As shown, the additional inference cost introduced by adaptive self-refinement leads to performance gains in both categories. Further removing model-aware data selection and the constructed self-refinement data both result in a clear performance drop, with the absence of model-aware data selection causing the most significant degradation. These findings demonstrate the effectiveness of our proposed modules, especially for the model-aware part which mainly comes from the novel definition of data sample difficulty.
4.3 Effects of Model-Aware Iterative Training
In this subsection, we explore several different choices when applying model-aware iterative training, to validate the effectiveness of our current proposal.
Choices of Selection Thresholds.
In our main experiments, we set $\alpha=0$ and $\beta=0.9$ to discard samples that are either extremely simple or too difficult for model training. Here we experiment with more different choices of the two thresholds. We turn to another three kinds of choices, preserving simpler samples, preserving more difficult samples, or preserving samples in more medium difficulty. Table 5 displays the evaluation results on BFCL. As shown, setting $\alpha=-1$ and $\beta=0.9$, which includes even extremely simple cases (i.e., $\alpha=0$), leads to a slight performance drop while increasing training time due to the inclusion of a larger training corpus. In contrast, preserving only relatively difficult cases – by setting $\alpha=0.1$ and $\beta=2$ – results in a notable performance degradation across both categories, with an overall decrease of approximately 1 point. This suggests that samples beyond the model’s current capability may hinder effective learning and limit performance improvements based on its existing potential. We further evaluate a stricter setting ($\alpha=0.1,\beta=0.8$), which discards more relatively simple and difficult cases. This also results in performance drop, especially for Live category. The likely reason is that an excessive number of training samples are discarded (reducing the dataset to approximately half the size of our main setting) while retaining only a small portion of the harder examples.
Thresholds Non-live Live Overall
$\alpha=0,\beta=0.9$ (ToolACE-R) 90.94 81.72 86.33
$\alpha=-1,\beta=0.9$ 90.73 81.27 86.00
$\alpha=0.1,\beta=2$ 89.27 81.35 85.31
$\alpha=0.1,\beta=0.8$ 90.83 80.38 85.61
Table 4: Accuracy performance on BFCL when applying different threshods during model-aware data selection. We always select samples that their difficulty score satisfy $\alpha<D<\beta$.
Models Non-live Live Overall
$k=8$ (ToolACE-R) 90.94 81.72 86.33
  + Adaptive SR 91.32 82.34 86.83
$k=1$ 90.94 80.98 85.96
  + Adaptive SR 90.71 82.01 86.36
Table 5: Accuracy performance on BFCL when setting $k$ value in Eq. 5 as 88 and 11. “SR” is short for “Self-Refine”.
Choices of $k$.
In Eq. 5, we define the difficulty score by allowing the model to make $k$ attempts, thereby better exploring its potential and producing a more robust estimate. To evaluate the impact of different $k$ values, we compare our main setting ($k=8$) with a reduced setting ($k=1$). The corresponding results are shown in Table 5. The model trained with $k=1$ exhibits a notable performance drop, particularly in the Live category, where difficult cases play a more critical role. Upon examining the final training set, we observe that approximately $10\%$ of the samples within the difficulty range $[0.8,0.9]$ under the $k=8$ setting are excluded when using $k=1$. This provides evidence that a larger $k$ value yields a more stable difficulty estimation and facilitates the selection of more appropriate training samples.
Choices of Models.
We further examine the “model-aware” part, where we use the base model to be trained to evaluate the difficulty of training samples. We conduct an additional experiment. Instead of using data selected by the base model Llama3.1-8B-Inst, we train Llama3.1-8B-Inst with data selected by Qwen2.5-3B-Inst and Qwen2.5-7B-Inst, as well as with the intersection of all three models’ selected data – i.e., only the samples that all three models perform similar (within the same difficulty scope) are included in the dataset. The results of the trained models on BFCL, along with the amounts of corresponding training data (in the first iteration), are presented in Table 6. As shown, performance drops when using data selected by other models, even when the data volume is larger. The intersection dataset, which contains less data, achieves comparable results. This observation underscores the benefits of model-aware data selection.
Models Live Overall Amount
ToolACE-R 81.72 86.33 120K
Select w/ Qwen2.5-3B 80.83 85.50 125K
Select w/ Qwen2.5-7B 81.27 85.97 134K
Intersection Data 81.37 86.08 108K
Table 6: Accuracy performance on BFCL when using different selected data to train Llama3.1-8B-Inst, along with the data amount. “Qwen2.5-3/7B” is short for Qwen2.5-3/7B-Inst.
Models Non-live Live Overall
Qwen2.5-7B-Inst 84.38 72.17 78.28
  + Original Data 88.29 78.39 83.34
  + ToolACE-R 89.23 81.20 85.21
Mistral-7B-Inst-v0.2 55.25 55.07 55.16
  + Original Data 84.06 70.17 77.12
  + ToolACE-R 87.98 80.01 84.00
Table 7: Accuracy performance of using different backbones on BFCL.
Iterative Training.
To examine the training dynamics, we visualize the iterative training process by illustrating the trends in both accuracy and the amount of training data across iterations, as shown in Figure 4. Iteration 0 corresponds to the base model, LLaMA3.1-8B-Inst. As depicted, the number of selected training samples decreases monotonically with each iteration, primarily because more samples are assigned a difficulty score of 0 and consequently excluded from the training corpus. Meanwhile, accuracy improves rapidly during the first two iterations, peaking at iterations 2 and 3. Beyond this point, performance plateaus, with no further gains observed in subsequent iterations. This pattern suggests that excessive training may lead to overfitting on the filtered dataset. Therefore, to save the training cost, we recommend iterative training at most 3 times.
4.4 Further analysis
Figure 3: Accuracy performance and training amount of each iteration during training. Iteration 0 means the model before training.
Figure 4: Accuracy Performance of Qwen2.5-Series. “Base” refers to the Qwen2.5-Instruct models, while “+ ToolACE-R” uses our method.
Different Backbones.
Model Size Scaling.
We further examine whether our method can be applied to models of varying sizes and whether performance improves as model size increases. For this, we use Qwen2.5-Instruct models of different sizes, including 0.5B, 1.5B, 3B, and 7B, as base models. The results, shown in Figure 4, display both the original performance and the performance after applying ToolACE-R. Applying ToolACE-R consistently improves performance across all sizes, particularly for smaller models like 0.5B and 1.5B. This demonstrates the effectiveness of ToolACE-R across a range of model sizes.
5 Conclusion
In this paper, we introduce ToolACE-R, a tool learning technique designed to enable large language models (LLMs) to effectively leverage external tools. ToolACE-R consists of two components: model-aware iterative training and adaptive self-refine iterative inference. Model-aware iterative training maximizes the model’s potential by iteratively selecting the most appropriate samples for training, with a novel definition of sample difficulty. Adaptive self-refine further enhances performance by allowing the model to autonomously determine when to stop refining its own answers during iterative inference. Experimental results demonstrate that ToolACE-R consistently improves performance across various tool-calling benchmarks and is effective with different backbone models. Additionally, our analyses highlight the impact of the proposed modules on model performance.
Appendix A Appendix
A.1 Algorithm for Adaptive Self-refine
Please refer to Alg. 1 for detailed process of adaptive self-refine strategy.
@@ A.1 Algorithm for Adaptive Self-refine
0: Model $\theta$, query $q$, tool candidates $T$, refine prompt $r$, maximum iteration time $n$
0: Refined final answer AA
[…]
@@
3:  $A_{i}=f_{\theta}(\langle q,T\rangle,A_{i-1},r)$  // Get the refined answer based on the answer of last iteration
@@
8:return  $A_{n}$  // Achieves the maximum iteration time
A.2 Detailed Introduction to Benchmarks